OpenStudy (anonymous):

The position of a particle at any time t>=0 is given by x(t)=t^2-2, y(t)=(2/5)t^3. Find the magnitude of the velocity vector at t=4.

OpenStudy (anonymous):

I know that the velocity vector is <2t, (6/5)t^2> but how do I find the magnitude of the velocity vector at t=4?

OpenStudy (anonymous):

I know the formula for finding the magnitude of the velocity vector at t=4 but the answer is 104/5. How do I simplify sqrt(4t^2+(36/25)t^4) and plug 4 in?

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

@ybarrap @kirbykirby

OpenStudy (kirbykirby):

\[\sqrt{4t^2+\left(\frac{36}{25}\right)t^4}=\sqrt{t^2\left(4+\left(\frac{36}{25}\right)t^2\right)}=t\sqrt{4+\left(\frac{36}{25}\right)t^2}=\\4\sqrt{4+\left(\frac{36}{25}\right)4^2}=4\sqrt{4+\left(\frac{36}{25}\right)16}=4\sqrt{\frac{100}{25}+\left(\frac{576}{25}\right)}=\\4\sqrt{\frac{676}{25}}=4\frac{\sqrt{4\cdot169}}{5}=4\frac{\sqrt{2^2\cdot13^2}}{5}=4\frac{2\cdot13}{5}=\frac{104}{5}\] Honestly I don't know what else you can do o.o I mean once you have the formula it is straightforward. But I'm assuming you are not allowed calculators? Because formula can't really be simplified much more the way it is...

OpenStudy (kirbykirby):

on the 2nd line I made the substitution t=4

OpenStudy (anonymous):

Is this the only way to solve this problem?

OpenStudy (kirbykirby):

Plug into a calculator :)?

OpenStudy (anonymous):

Thanks.

OpenStudy (kirbykirby):

But yeah really there is not much that can be done here... I mean the hard part of the problem was solved by you! Finding 104/5 can be done by calculator, or by hand calculation above... but if you have access to the calculator I highly recommend using it as you save a lot of time.